Reports 1-1 of 1 Clear search Modify search
Environmental Monitoring (Environmental Monitoring)
Paoletti, Fiori, Tringali - 16:33 Friday 24 June 2022 (56249) Print this report
Temporary Magnetic sensor added near the CEB HALL AHU and preliminary shutdown

Today starting from 08:00 UTC we've added a magnetic probe in the CEB HALL AHU room (see photo).

Its name is ENV_CEB_AHU_MAG

In the FFT plot we can see some relevant lines,below is a list:

Hz Origin
3.74 Return fan
5.92 unknown
7.46 unknown
7.65 Supply fan
11.2 unknown
11.4 unknown
15.3 unknown
18.5 unknown
45 Return inverter
46 Supply inverter
47.77 unknown (50Hz lower sideband)
52.23 unknown (50Hz higher sideband)

Note that the sidebands spacing (2.23 Hz) is the same as the Supply Belt rate frequency. Maybe just a coincidence.

Most (if not all) of the known lines have high coherence with many seismometers and accelerometers.

A quick shutdown of the AHU (few minutes) has been performed around 13:00 UTC, below is the timeline:

13:02 Return OFF

13:07 Supply OFF

13:13 Return ON

13:13:30 Supply ON

It seems that the 50 Hz sidebands are associated to the Supply system

The Central Hall temperature raised about 0.6 degrees (from 21.5 to 22.1)

Images attached to this report
Comments to this report:
Paoletti - 9:17 Monday 27 June 2022 (56269) Print this report
Images attached to this comment
fiori, tringali - 10:18 Friday 05 August 2022 (56657) Print this report

The simple exercise in Figure below would demonstrate that the 50 Hz sidebands at +- 2.23 Hz seen in the magnetometer next to the CEB AHU are indeed associated to the supply fan belts.

The signal of the accelerometer placed on the supply fan is low pass filtered (4 Hz, for convenience) and multiplied by one CEB IPS voltage monitor (measuring essentially the CEB 50Hz signal).

Figure 1 shows (top plot) the coherence between the signal resulting from the operation described above and the magnetometer placed close to the CEB AHU. There is a significant coherence (around 0.8) at the location of the +- 2.23 Hz sidebands. The bottom plot is the ASD of the magnetometer.

Images attached to this comment
fiori, paoletti, tringali - 18:20 Friday 05 August 2022 (56667) Print this report

Today we moved the magnetic probe ENV_CEB_AHU_MAG close to the main electric switch board and the CEB AHU inverters. The new location is shown in the pictures. The aquired channel is the "Z" which is oriented along the longer side of the probe. Analysis will follow.

Images attached to this comment
Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×